12 Pablos utility function is U(x,y)= x+10y- y2/2, where x is the number of xs he consumes per week and y is the number ofhow do you solve without lagrangians?
Answer #1

Initial MRS=-MU/MUy= -1/(10-y)

Slope of Budget constraint= -Px/Py= -1/5

optimal bundle:

-1/5=-1/(10-y)

10-y=5

y*= 5

Budget constraint: x+5y=200

x=200-5*5=175

x*=175

U= 175+10*5-5^2/2=175+50-12.5= 212.5

When Py=1

slope of budget constraint= -1/1=-1

-1=-1/(10-y)

10-y=1

y*= 9

x+y=200

x*=200-9= 192

U= 191+10*9-9^2/2= 240.5

The maximum price Pablo will be willing to pay= 240.5-212.5=$28

So, Answer=$28